### v1

# Extreme value distribution for GRB prompt data

Abstract
Context. Gamma-Ray Bursts (GRBs) are known to be unpredictable in time and position. A few (observationally) exceptional events have been observed, as GRB 221009A that stands out for its fluence and peak flux, being orders of magnitude higher than what measured so far.
Aims. Analyzing the observed fluence, peak flux or duration distributions typically requires one to assume some scenarios, and the consistency of the observed data with the predictions turns out to be an important model diagnostic. However, it is also of interest to model these distributions using general statistical properties that do not rely on specific model assumptions, allowing one to derive inferences only based on the consistency of the observed distributions with the hypothesis of one single population of events that generate them.
Methods. We obtained fluences, peak fluxes and durations from the catalogues of GRBs observed by the CGRO-BATSE and Fermi-GBM instruments. We selected the extreme values in slots of equal duration and modelled their distributions by the generalized extreme value (GEV) formalism. The GEV distribution is a limit distribution naturally arising when the number of observations is large and is essentially independent of the phenomena producing the observed data.
Results. The distributions of extreme values for fluences, peak fluxes and durations are consistent with being extracted from a single population of events but the fluence and peak flux recorded for GRB 221009A constitutes a striking exception. The probability to observe such an event, assuming it is a cosmological GRB, is low, with a median value of about one event per millennium for the fluence and about one event per century for the peak flux.
Key Words.:
Gamma-ray burst: general – Gamma-ray burst: individual: GRB 221009A – Methods: statistical1 Introduction
Gamma-Ray Bursts (GRBs) have been routinely observed by high-energy instrument since their discovery by the Vela early warning satellites (Klebesadel et al., 1973). Since then, a fairly large set of spaceborne instruments have monitored the sky at times, generating catalogues with thousands of events, such as those reported by the Burst and Transient Source Experiment (CGRO-BATSE, Goldstein et al., 2013)111https://heasarc.gsfc.nasa.gov/W3Browse/cgro/bat5bgrbsp.html, Konus-Wind (e.g., Tsvetkova et al., 2017), and Fermi Gamma-Burst Monitor (Fermi-GBM, Gruber et al., 2014; von Kienlin et al., 2014; Bhat et al., 2016; von Kienlin et al., 2020; Poolakkil et al., 2021)222https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermigbrst.html.
GRBs show two usually distinct emission phases, the prompt and the afterglow (Zhang, 2018). They have been classified in (at least) two different flavors, the long soft and short hard GRBs (Kouveliotou et al., 1993; Qin et al., 2013), originating from two different classes of progenitors (Zhang, 2018). However, exceptions to a classification based only on duration and hardness ratios have been singled out (see, e.g., Ahumada et al., 2021; Zhang et al., 2021; Yang et al., 2022; Dichiara et al., 2023; Levan et al., 2024; Yang et al., 2024; Yi et al., 2025).
The availability of such a large collection of data has naturally allowed for a large variety of statistical analysis about, e.g., their sky distribution (Briggs et al., 1996), their duration and hardness (Kouveliotou et al., 1993), or their luminosity function (Fenimore et al., 1993; Salvaterra et al., 2012) or relations between prompt and/or afterglow parameters (Yonetoku et al., 2004; Amati, 2006; Ghirlanda et al., 2012; Dainotti et al., 2023). In this framework, the rather recent detection of GRB 221009, with its extreme properties in terms of fluence and peak flux (see, e.g., Frederiks et al., 2023; Burns et al., 2023; Lesage et al., 2023), represents a formidable test bench for any attempt to model the observed population(s) of cosmological GRBs and even the possibility of physics beyond the standard model (Galanti et al., 2024, 2025).
GRB 221009A (e.g., Malesani et al., 2023, and references therein), was rapidly dubbed the “BOAT”, the “Brightest Of All Times”. Indeed, it was an exceptionally bright GRBs producing a fluence and peak flux (after correction for saturation) more than an order of magnitude greater than any other event detected so far (Burns et al., 2023, and references therein). This event is of course interesting by itself, but several studies have also been devoted to establish the probability of observing such an event if GRB 221009A were a regular cosmological GRBs. The results unavoidably depend on the various assumptions involved in the analysis. In Malesani et al. (2023) rates range from one case every few decades up to a century. Instead, Burns et al. (2023) obtained, by geometric extrapolation of the distributions for a multi-mission catalogue of long GRBs, that the fluence of GRB 221009A makes it a once-in-10,000-year event. Estimates in 1 for to years have been obtained by several more authors (e.g., Atteia et al., 2025; Lan et al., 2023; O’Connor et al., 2023; Williams et al., 2023). Such a low rate stimulated the researchers to find different possibilities to avoid the (statistical) contradiction between the actual observation and the length of the sky monitoring (a few decades). Bloom (2022) suggested that GRB 221009A could be the result of Galactic lensing. Finke & Razzaque (2024) discussed the possibility that GRB 221009A belonged to a different population of narrow-beam GRBs, so that its occurrence rate could be as high as approximately one per 200 yr. Eventually, the possibility that this event was not a cosmological GRB has also been discussed (e.g., Navia et al., 2024).
The problem of determining the true rate of events comparable to GRB 221009A is therefore still open and full of consequences for our understanding of the cosmological GRB population. In this paper, we propose a different approach, almost model-independent, i.e., based only on the assumptions that the detection probability of bright GRBs by high-energy satellites is essentially constant during the lifetime of the mission and that the probability of occurrence of such an event is also constant in time, i.e., it is stationary. In fact, we modelled the set of bright events observed by the CGRO-BATSE and Fermi-GBM catalogues by the so-called generalized extreme value distribution (GEV, Beirlant et al., 2004; Coles, 2001), i.e. a statistical distribution naturally emerging asymptotically and almost independently of the phenomena under study. In Sect. 2 we describe the properties of the GEV distribution, in Sect. 3 we introduce the data set used in this work and in Sect. 4 the analysis carried out. Finally, in Section 5 we comment our results and present our conclusions.
2 The Generalized Extreme Value distribution
The extreme value theory, i.e. the attempt to evaluate the probability of occurrence of extreme measurements, is a rapidly evolving field of modern statistics. It is applied to many scenarios, i.e., climatology, engineering, physics, and economy (Coles, 2001). In astronomy, recent applications include the modelling of the brightest galaxies in a survey (Heather et al., 2024), the number of sunspots (Zhang et al., 2024), the maxima of periodograms in time-series analysis (Gondhalekar et al., 2023; Süveges, 2014), density fields in cosmology (Repp & Szapudi, 2018), masses of galaxy clusters (Waizmann et al., 2012), etc.
The most interesting aspect of the GEV distribution is that it can be proved that it holds asymptotically essentially independently of the probability distribution describing the phenomena that generate the observed data. The mathematical proof is somehow tedious (e.g., Beirlant et al., 2004) but not difficult, and there is some analogy with the central limit theorem that guarantees that, under rather general conditions, the limiting distribution of sample means for a large family of parental statistical distributions is a standard normal distribution (Fischer, 2011).
If we collect the maxima, , of a large number of random variables, , identically distributed according to a given continuous distribution, , with , the GeV distribution can be written as:
where .
The parameter is usually called shape and it is related to the tail decay of the underlying distribution of . is the mean value of the distribution and is the scale. The shape divides the GEV family into three subfamilies. Negative shape parameters signal distributions of maxima of variables with a finite upper boundary (i.e., , while a positive or zero shape parameter indicate that there is no upper limit. The case is also known as the “Gumbel” distribution. It is defined as the limit function when , and takes the form:
with .
A GEV model is trained typically dividing a dataset into blocks of uniform length, e.g. a year, etc. Then, from each block of data, the extreme (maximum or minimum) is extracted, and finally a GEV model is fitted to the set of extreme values. Once trained, the GEV distribution can be used to calculate return levels. The return level associated with the return period , is simply the level expected to be exceeded once in all blocks of the same length . Credible intervals for any estimated quantity can be derived by the posterior distribution of the GEV parameters.
The choice of the length of the blocks used to extract extreme values is not driven by first principles, and there is a well-known bias-variance trade-off. The GEV is a limit distribution, therefore longer blocks, i.e. including a large number of elements, provide a better approximation. On the other hand, longer blocks, for any finite set of observed data, imply a reduced number of elements and unavoidably a poorer prediction with a larger variance. Smaller blocks might, on the contrary, give a worse approximation of a GEV distribution, yielding biased predictions. Therefore, the choice of block length, unless driven by physical reasons, is a trial and test process, and the results will be tested using appropriate model diagnostics as, in most cases, the quantile-quantile (QQ) plot (e.g., Coles, 2001). If the model is good, the points will closely follow a straight line with intersect 0 and slope 1 without strong systematic deviations.
We mention, finally, that a crucial assumption of extreme value analysis is that the original distribution of data where extreme values are extracted from should be stationary. If this does not happen it is indeed possible to include in the analysis exogenous variables to model the non-stationarity (e.g. trends, seasonality, etc., Coles, 2001).
3 Data set selection
In this work, we based our analyses on the Fermi-GBM GRB catalogue (e.g., Poolakkil et al., 2021) and we also considered the CGRO-BATSE GRB catalogue (Goldstein et al., 2013), as a comparison case with no events comparable to GRB 221009A. The GBM catalogue contains 4029 events from July 2008 to June 2025333Update up to June 26, 2025. , while the BATSE catalogue contains 2145 events from April 1991 to August 1996. From these catalogues, we extracted the trigger time, the fluence, the peak flux and the GRB duration, expressed, as usual, by the parameter , the duration of the time interval during which 90% of the total observed counts were detected. In more detail, from the BATSE catalogue we considered the total fluence in the energy range 20-300 keV (summing the fluences recorded in the 20-50, 50-100 and 100-300 keV bands), in units of erg cm-2 and the peak flux in the 20-300 keV energy range on the 1024 ms timescale in units of photons cm-2 sec-1. From the GBM catalogue we considered the fluence in the 10-1000 keV energy range, in erg and the peak flux in the same energy range on the 1024 ms timescale in units of photons cm-2 sec-1. Choosing different baselines for the peak flux computation does not change any of the conclusions reported in this paper.
Since we are interested in the extreme values for the considered parameters, it is necessary to discuss the various biases affecting data obtained from the selected catalogues. In fact, the brightest events, are often affected by saturation. A correction for the saturation is typically carried out comparing results from different instruments by a suitable cross-calibration, and/or a modelling of the spectral evolution of the event in order to infer the expected behaviour when the considered instrument was unable to provide reliable information. In most cases, such a correction is of limited entity and not important for the present study. GRB 221009A is an exception due to its peculiar brightness. A saturation-corrected fluence was computed in Lesage et al. (2023) and the reported value is more than two times higher than the value reported in the GBM catalogue. A saturation corrected peak-flux was computed by Frederiks et al. (2023), basing on Konus-Wind data in an energy band close to the Fermi-GBM one. This estimate turns out to be an order of magnitude higher than what reported in the GBM catalogue. In the following analysis we adopted the saturation-corrected values for the BOAT.
In addition, the BATSE catalogue is known to be incomplete toward the highest fluxes and fluences (see, e.g., Kaneko et al., 2006; Burns et al., 2023), due to several factors, as saturation and triggering strategy. This unavoidably affects our capability to derive a solid GEV distribution for the parameter of our interest and incompleteness in the derived extreme values can make more uncertain the extrapolation toward the highest fluences and peak fluxes, beyond the statistical uncertainties. On the other hand, as far as the bulk of the high fluence and flux reported events are reliable, the global features of the computed GEV distributions should be still of interest.
Independently of the considered catalogue, events with very long duration might suffer from biases due to various possible constrains affecting instrument efficiency aboard of low orbit satellites (e.g. Earth occultation, South-Atlantic anomaly crossing, telemetry limitations, etc.). Catalogues are therefore likely truncated missing or incorrectly recording some of longest durations GRBs. The BATSE catalogue does not report events longer than s, while the GBM catalogue seems to be limited at , indicating the higher level of incompleteness suffered from the former. A few ultra-long duration GRBs () have also been singled out (e.g., Levan et al., 2014), opening a long-lasting debate about their possible different origin from the bulk of cosmological GRBs. Deriving the true duration of a very long GRB can be a difficult task, often relying on data from multiple missions. Inclusion of these data in the present analysis could undoubtedly be of interest, but would also need a proper homogenization since we should derive a duration seen by the given (BATSE or GBM) instrument without limitations, i.e., a different quantity with respect to the duration of emission activity of the source at any energy. Since the focus of the present work is on the possible peculiarity of GRB 221009A, whose duration is well within the recorded values in the GBM catalogue, we decided not to apply any correction to the duration data.
4 GEV analysis
The choice of the block length to extract the maxima, as discussed in Sect. 2, is not driven by first principles and is a trade-off between the goodness of the GEV distribution fit vs. the variance of the prediction. The higher the number of events per block (i.e., the longer the block duration) the better is the GEV distribution fit and the worse is the variance on the prediction. In our cases, block length of 30 days seems to offer a good compromise given the average number of events per day in both catalogues ( day). For the BATSE catalogue we obtained a sample of 66 extreme values. The richer GBM catalogue instead allowed us to derive a sample 207 extreme values. The selected ”extreme” events are shown in Fig. 3.
For both the considered catalogues, the selected extreme values have a signal-to-noise ratio better than , making their uncertainties irrelevant for the considerations derived in this work.
We performed a Bayesian analysis and adopted improper uniform priors for the model parameters. The resulting posterior is proper as long as the sample size is larger than 3 (Northrop & Attalides, 2016). Sampling is carried out by the “No-U-Turn Sampler” extension (Hoffman & Gelman, 2014) to the Hamiltonian Monte Carlo algorithm (Brooks et al., 2011). The sampled chains in all cases converge to equilibrium and the posterior distribution of parameters shows a peaked distribution for the three parameters of the model. Corner plots for the six cases considered in this work are shown in Fig. 4 and summary statistics for the GEV distribution fitted parameters are reported in Table 1.
In several cases the shape parameter distributions include the zero values, the simple Gümbel distribution, in their possible ranges. However, some of the distributions seem to be better fit by negative shape parameters, although in no case the evidence is very strong. This essentially implies that the available datasets are not able to unambiguously indicate whether their extreme value distributions are characterized by an asymptotic maximum value or are likely ever increasing going to an infinite number of blocks. There are also instrumental limitations to consider in evaluating the obtained results. E.g., the difficulties in recording long uninterrupted observations makes the lack of very long duration GRBs in the considered catalogues, at least partly, artificial, likely pushing the shape parameter to negative values.
Since the fluence and peak flux of GRB 221009A are highly discrepant with respect to the population of GRBs from the GBM catalogue (see also Sect. 5), we also carried out the fit excluding the BOAT from the dataset. This is, at this stage, a rather arbitrary choice but, quite interestingly, without GRB 221009A the GEV distribution for the fluence would indicate an upper boundary value at erg . This figure is close to the fluence recorded for the BOAT. The analogous limit for the peak flux is instead much higher than the recorded value and, as expected, the GEV distribution for the duration is almost unaffected by the inclusion of GRB 221009A in the analysis.
As discussed in Sect. 2, the fit quality was checked by diagnostic plots, as the widely known QQ plots. For the both the datasets considered in this work, the distributions of durations, peak fluxes and fluences turn out to be nicely fit by a GEV distribution (Fig. 1), with the already mentioned notable exception of GRB 220109A for the fluence and the peak flux that stands clearly out among the extreme values from the GBM catalogue. As already pointed out in past analyses (e.g., Burns et al., 2023; Malesani et al., 2023), the peculiarity of GRB 220109A is due to a combination of factors, i.e., persistent high flux emission, long duration and the spectral hardness.
Finally, the return plots are shown in Fig.2. The reported levels are expected to be exceeded on average once every years, i.e., the return level is exceeded in any particular block with probability .
For the BATSE catalogue, as expected from the analysis of the QQ plot, the return levels are all well consistent with the modelling (Fig. 2). Things are considerably different for the GBM catalogue data, where again GRB 220109A appears grossly inconsistent with the model.
Just extrapolating the fit distribution, the median value of the model return level, at the fluence of GRB 220109A, is more than years, and the BOAT fluence enters the 95% allowed range at years. A peak flux value comparable to what shown by GRB 220109A is expected in years, while the observed peak flux enters the 95% range in years. Removing GRB 220109A the GEV distributions for the fluence and peak flux give a negligible probability of observing an event of comparable fluence or peak flux for time scales of thousands years.
5 Discussion and Conclusions
The most interesting feature of a GEV analysis is that it does not require any assumptions beyond that the analysed data are drawn from a stationary distribution. Durations, peak fluxes and fluences for the brightest events detected by a specific instrument are satisfying this requirements, since, to a large extent, we can assume that the natural degradation of the instruments have a minor or even negligible impact on their capability to detect observationally bright events. In addition, focusing on the longest, brightest and with the highest fluence events, we are also automatically excluding from the data sets GRBs belonging to the short duration class (Kouveliotou et al., 1993) with the partial exception of the extreme values for the peak fluxes. Quite interestingly, for both the considered catalogues, a percentage of 14% (BATSE) and (GBM) of peak flux extreme values is due to events shorter than the canonical 2 s value. On the other hand, in spite of the fact that the brightest events are also, unavoidably, the best studied, there is still the (more or less virtual) possibility that the BATSE and GBM samples are polluted by very bright events due to phenomena erroneously classified as GRBs. As a matter of fact, it was also proposed that some of the brightest events come from a different (and subdominant) population of GRBs (e.g., Finke & Razzaque, 2024; Briel et al., 2025). Analogous considerations have been proposed for the ultra-long GRBs ( 1000 s) (e.g., Levan et al., 2014) that, however, are not included in the present analysis.
Our results indicate that, apart from the case of GRB 221009A, that we are going to discuss later, the GEV distributions for GRB reported in the BATSE and GBM catalogues are consistent with being drawn from a single population of events. Their respective GEV distributions are also substantially consistent with each other (Table 1). For instance, the return level for a GRB duration of s is close to one event in 20-30 years in both cases, possibly inconsistent with the few ultra-long GRBs observed so far, although the inefficient recording of very long duration events makes any conclusions about this topic at this stage highly preliminary.
As already observed, the data for GRB 221009A are well consistent with the population GEV distributions if we refer to the duration. They are instead highly discrepant for the fluence and peak flux. This is a non trivial point, and shows that the peculiarities of the BOAT are due to the combination of several factors as, long-lasting high-flux emission (and its spectral hardness, e.g., Frederiks et al., 2023) rather than to a very long or very bright flux only. This is also in agreement with findings from previous works (e.g., Malesani et al., 2023; Burns et al., 2023). The conclusions about the peak flux hold even if different baselines for the flux computations are used.
The fluence and peak flux of GRB 221009A are, therefore, grossly inconsistent with the GEV distribution. The median of the return level for the fluence yields a probability of having one event of this fluence in about 1000 years, while for the peak flux the median return value is about 140 years. All in all, GRB 221009A appears an unlikely event for its fluence and peak flux, given the duration of the monitoring of the considered missions. If we, on the contrary, adopt the hypothesis that an event with fluence and peak flux as large as or larger than GRB 221009A is unlikely, this might imply that one of the assumptions of the present analysis, i.e., that the measured extreme values are drawn from an homogenous population of GRBs, is violated. As a matter of fact, the possibility that GRB 221009A is not a typical cosmological GRB has been proposed and discussed in the literature. This might be due to a different geometry of the event (Finke & Razzaque, 2024), a misclassification of the event as a GRB (Navia et al., 2024) or just low-probability intense gravitational lens enhancement (Bloom, 2022). In all cases, these hypotheses require that their rate of occurrence is low enough not to affect the bulk of the GRB population observations and that they affect mainly the fluence, and to a lower extent the peak flux, but not duration of the detected events.
An analysis considering all the GRBs detected by any instrument after a proper homogenization is in principle possible, although the possibility of introducing artefacts due to the incomplete knowledge of emission spectra and the intrinsic limitations of the involved instruments is high. We leave this possible development to a future work.
6 Data availability
We developed software tools and used third-party libraries coded in the julia language (v. 1.13.4-1.13.5, Bezanson et al., 2017)444https://julialang.org/. GEV models and fitting tools are from the Extremes package555https://jojal5.github.io/Extremes.jl/dev/ (v. 1.0.3-1.0.5, Jalbert et al., 2024). Plots are produced by the makie666https://docs.makie.org/v0.22/ (v. 0.22.2-0.24.0, Danisch & Krumbiegel, 2021) with the CairoMakie backend, and corner plot are produces by the PairPlots777https://sefffal.github.io/PairPlots.jl/dev/ (v. 3.0.1, Thompson, 2023) packages. Project management was carried out by the DrWatson888https://juliadynamics.github.io/DrWatson.jl/dev/ package (v. 2.18.0, Datseris et al., 2020). Bibliographic research has also been aided by the Google Gemini (v. 2.5) IA model.999https://gemini.google.com/
Data and code used in this work are available at the gitlab repository: https://www.ict.inaf.it/gitlab/stefano.covino/grb-extreme-values.git.
Acknowledgements.
I thank dr. Giancarlo Ghirlanda and Om Sharan Salafia for useful discussions and dr. Eric Burns for kindly sharing his datasets. I also thank the anonymous referee. Her/his pertinent comments have substantially improved this manuscript. I also acknowledge the contribution by the Italian Space Agency, contract ASI/INAF n. I/004/11/6.References
- Ahumada et al. (2021) Ahumada, T., Singer, L. P., Anand, S., et al. 2021, Nature Astronomy, 5, 917
- Amati (2006) Amati, L. 2006, MNRAS, 372, 233
- Atteia et al. (2025) Atteia, J.-L., Bouchet, L., Dezalay, J.-P., et al. 2025, ApJ, 980, 241
- Beirlant et al. (2004) Beirlant, J., Goegebeur, Y., Segers, J., & Teugels, J. 2004, Statistics of Extremes: Theory and Applications (Wiley), pagination: 522
- Bezanson et al. (2017) Bezanson, J., Edelman, A., Karpinski, S., & Shah, V. B. 2017, SIAM Review, 59, 65
- Bhat et al. (2016) Bhat, P. N., Meegan, C. A., von Kienlin, A., et al. 2016, VizieR Online Data Catalog: The third Fermi/GBM GRB catalog (6yr) (Bhat+, 2016), VizieR On-line Data Catalog: J/ApJS/223/28. Originally published in: 2016ApJS..223…28B
- Bloom (2022) Bloom, J. S. 2022, Research Notes of the American Astronomical Society, 6, 220
- Briel et al. (2025) Briel, M. M., Fragos, T., Salafia, O. S., et al. 2025, arXiv e-prints, arXiv:2502.09187
- Briggs et al. (1996) Briggs, M. S., Paciesas, W. S., Pendleton, G. N., et al. 1996, ApJ, 459, 40
- Brooks et al. (2011) Brooks, S., Gelman, A., Jones, G., & Meng, X.-L. 2011, Handbook of Markov chain Monte Carlo (CRC press)
- Burns et al. (2023) Burns, E., Svinkin, D., Fenimore, E., et al. 2023, ApJ, 946, L31
- Coles (2001) Coles, S. 2001, An introduction to statistical modeling of extreme values, Springer Series in Statistics (London: Springer-Verlag)
- Dainotti et al. (2023) Dainotti, M. G., Lenart, A. Ł., Chraya, A., et al. 2023, MNRAS, 518, 2201
- Danisch & Krumbiegel (2021) Danisch, S. & Krumbiegel, J. 2021, Journal of Open Source Software, 6, 3349
- Datseris et al. (2020) Datseris, G., Isensee, J., Pech, S., & Gál, T. 2020, Journal of Open Source Software, 5, 2673
- Dichiara et al. (2023) Dichiara, S., Tsang, D., Troja, E., et al. 2023, ApJ, 954, L29
- Fenimore et al. (1993) Fenimore, E. E., Epstein, R. I., Ho, C., et al. 1993, Nature, 366, 40
- Finke & Razzaque (2024) Finke, J. D. & Razzaque, S. 2024, ApJ, 975, 70
- Fischer (2011) Fischer, H. 2011, A History of the Central Limit Theorem: From Classical to Modern Probability Theory (New York, NY: Springer New York)
- Frederiks et al. (2023) Frederiks, D., Svinkin, D., Lysenko, A. L., et al. 2023, ApJ, 949, L7
- Galanti et al. (2024) Galanti, G., Nava, L., Roncadelli, M., Tavecchio, F., & Bonnoli, G. 2024, arXiv e-prints, arXiv:2412.21175
- Galanti et al. (2025) Galanti, G., Roncadelli, M., Bonnoli, G., Nava, L., & Tavecchio, F. 2025, arXiv e-prints, arXiv:2502.03453
- Ghirlanda et al. (2012) Ghirlanda, G., Nava, L., Ghisellini, G., et al. 2012, MNRAS, 420, 483
- Goldstein et al. (2013) Goldstein, A., Preece, R. D., Mallozzi, R. S., et al. 2013, ApJS, 208, 21
- Gondhalekar et al. (2023) Gondhalekar, Y., Feigelson, E. D., Caceres, G. A., Montalto, M., & Saha, S. 2023, ApJ, 959, L16
- Gruber et al. (2014) Gruber, D., Goldstein, A., Weller von Ahlefeld, V., et al. 2014, ApJS, 211, 12
- Heather et al. (2024) Heather, C., Chantavat, T., Chongchitnan, S., & Silk, J. 2024, MNRAS, 534, 173
- Hoffman & Gelman (2014) Hoffman, M. D. & Gelman, A. 2014, Journal of Machine Learning Research, 15, 1593
- Jalbert et al. (2024) Jalbert, J., Farmer, M., Gobeil, G., & Roy, P. 2024, Journal of Statistical Software, 109, 1–35
- Kaneko et al. (2006) Kaneko, Y., Preece, R. D., Briggs, M. S., et al. 2006, ApJS, 166, 298
- Klebesadel et al. (1973) Klebesadel, R. W., Strong, I. B., & Olson, R. A. 1973, ApJ, 182, L85
- Kouveliotou et al. (1993) Kouveliotou, C., Meegan, C. A., Fishman, G. J., et al. 1993, ApJ, 413, L101
- Lan et al. (2023) Lan, L., Gao, H., Li, A., et al. 2023, ApJ, 949, L4
- Lesage et al. (2023) Lesage, S., Veres, P., Briggs, M. S., et al. 2023, ApJ, 952, L42
- Levan et al. (2024) Levan, A. J., Gompertz, B. P., Salafia, O. S., et al. 2024, Nature, 626, 737
- Levan et al. (2014) Levan, A. J., Tanvir, N. R., Starling, R. L. C., et al. 2014, ApJ, 781, 13
- Malesani et al. (2023) Malesani, D. B., Levan, A. J., Izzo, L., et al. 2023, arXiv e-prints, arXiv:2302.07891
- Navia et al. (2024) Navia, C., Oliveira, M., Felicio, B., & Nepomuceno, A. 2024, arXiv e-prints, arXiv:2410.18131
- Northrop & Attalides (2016) Northrop, P. J. & Attalides, N. 2016, Statistica Sinica, 26, 721
- O’Connor et al. (2023) O’Connor, B., Troja, E., Ryan, G., et al. 2023, Science Advances, 9, eadi1405
- Poolakkil et al. (2021) Poolakkil, S., Preece, R., Fletcher, C., et al. 2021, ApJ, 913, 60
- Qin et al. (2013) Qin, Y., Liang, E.-W., Liang, Y.-F., et al. 2013, ApJ, 763, 15
- Repp & Szapudi (2018) Repp, A. & Szapudi, I. 2018, MNRAS, 473, 3598
- Salvaterra et al. (2012) Salvaterra, R., Campana, S., Vergani, S. D., et al. 2012, ApJ, 749, 68
- Süveges (2014) Süveges, M. 2014, MNRAS, 440, 2099
- Thompson (2023) Thompson, W. 2023, PairPlots.jl Beautiful and flexible visualizations of high dimensional data, https://sefffal.github.io/PairPlots.jl/dev
- Tsvetkova et al. (2017) Tsvetkova, A., Frederiks, D., Golenetskii, S., et al. 2017, ApJ, 850, 161
- von Kienlin et al. (2014) von Kienlin, A., Meegan, C. A., Paciesas, W. S., et al. 2014, ApJS, 211, 13
- von Kienlin et al. (2020) von Kienlin, A., Meegan, C. A., Paciesas, W. S., et al. 2020, ApJ, 893, 46
- Waizmann et al. (2012) Waizmann, J. C., Ettori, S., & Moscardini, L. 2012, MNRAS, 422, 3554
- Williams et al. (2023) Williams, M. A., Kennea, J. A., Dichiara, S., et al. 2023, ApJ, 946, L24
- Yang et al. (2022) Yang, J., Ai, S., Zhang, B.-B., et al. 2022, Nature, 612, 232
- Yang et al. (2024) Yang, Y.-H., Troja, E., O’Connor, B., et al. 2024, Nature, 626, 742
- Yi et al. (2025) Yi, S. X., Wang, C. W., Shao, X., et al. 2025, ApJ, 985, 239
- Yonetoku et al. (2004) Yonetoku, D., Murakami, T., Nakamura, T., et al. 2004, ApJ, 609, 935
- Zhang (2018) Zhang, B. 2018, The Physics of Gamma-Ray Bursts (Cambridge University Press)
- Zhang et al. (2021) Zhang, B. B., Liu, Z. K., Peng, Z. K., et al. 2021, Nature Astronomy, 5, 911
- Zhang et al. (2024) Zhang, R., Chen, Y.-Q., Zeng, S.-G., et al. 2024, Journal of Astrophysics and Astronomy, 45, 14
Appendix A Extreme values
We show here (Fig. 3) the extreme values for fluences, peak fluxes, and durations for GRB data reported in the CGRO-BATSE and Fermi-GBM catalogues in bins of 30 days.
Appendix B Corner plots and fitted values
We show here (Fig. 4) the ”corner plots” derived by the Bayesian analysis of extreme values for fluences, peak fluxes, and durations for GRB data reported in the CGRO-BATSE and Fermi-GBM catalogues. Summary statistics for the fitted parameters are shown in Table 1. Errors on the credible intervals are reported at 1 level.
Fluence | CGRO-BATSE | Fermi-GBM | Fermi-GBM |
(without the BOAT) | |||
Peak flux | CGRO-BATSE | Fermi-GBM | Fermi-GBM |
(without the BOAT) | |||
Duration | CGRO-BATSE | Fermi-GBM | Fermi-GBM |
(without the BOAT) | |||